### v1

# Optimizing Treatment Allocation in the Presence of Interference

Abstract
In Influence Maximization (IM), the objective is to – given a budget – select the optimal set of entities in a network to target with a treatment so as to maximize the total effect. For instance, in marketing, the objective is to target the set of customers that maximizes the total response rate, resulting from both direct treatment effects on targeted customers and indirect, spillover, effects that follow from targeting these customers. Recently, new methods to estimate treatment effects in the presence of network interference have been proposed. However, the issue of how to leverage these models to make better treatment allocation decisions has been largely overlooked. Traditionally, in Uplift Modeling (UM), entities are ranked according to estimated treatment effect, and the top entities are allocated treatment. Since, in a network context, entities influence each other, the UM ranking approach will be suboptimal. The problem of finding the optimal treatment allocation in a network setting is combinatorial and generally has to be solved heuristically. To fill the gap between IM and UM, we propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference to find solutions to the IM problem using treatment effect estimates. OTAPI consists of two steps. First, a causal estimator is trained to predict treatment effects in a network setting. Second, this estimator is leveraged to identify an optimal treatment allocation by integrating it into classic IM algorithms. We demonstrate that this novel method outperforms classic IM and UM approaches on both synthetic and semi-synthetic datasets.
Keywords Analytics  $\cdot$ Causal Inference  $\cdot$ Uplift Modeling  $\cdot$ Influence Maximization  $\cdot$ Network Analysis
1 Introduction
A well-studied and important problem within network analysis is Influence Maximization (IM) (Kempe et al., 2003; Domingos & Richardson, 2001; Richardson & Domingos, 2002; Azaouzi et al., 2021). The primary goal in IM is to identify the set of entities in a network to target with a particular intervention, e.g., a vaccine or a marketing campaign, in order to maximize the intervention’s effect across the entire network given a budget $k$. The effect of treating one entity can cascade through the network, affecting connected entities as well. Consequently, careful analysis is essential to allocate treatment efficiently. For instance, in viral marketing, (potential) customers influence each other’s purchasing decisions (Goldenberg et al., 2001). Hence, when determining whom to target with a marketing campaign, merely estimating the Individual Treatment Effects (ITEs), as commonly done in Uplift Modeling (UM) (Devriendt et al., 2018), may not suffice. Considering interference effects – where treating one entity may influence others – is critical to identify the optimal set of entities to target. Practical applications of this problem include, e.g., vaccination and marketing campaigns.
Influence Maximization.  The IM problem can be formalized as follows: given a network $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges, the objective is to identify the set of $k$ seed nodes that are to be treated in order to maximize the total influence spread, i.e., the total effect of treatment allocation (Kempe et al., 2003). In the IM literature, the influence spread of a certain treatment allocation is typically modeled using a diffusion process. This process describes how information or behavior disseminates through the network (Granovetter, 1978; Goldenberg et al., 2001).
We identify three significant limitations in the current approaches to address the IM problem. Firstly, node features are mostly ignored. Nevertheless, these features may contain important information regarding the spread of behaviors or effects through the network (Ogburn & VanderWeele, 2014; Ma & Tresp, 2021). Secondly, to the best of our knowledge, existing methods typically assume a homogeneous effect of treatment. More specifically, it is assumed that all entities will respond positively (e.g., purchase the product) if they receive treatment. Yet, in many applications, treatment effects vary among entities, and the susceptibility to treatment often depends on entity characteristics (Shalit et al., 2017; Athey & Imbens, 2016). Lastly, the most commonly used diffusion processes rely on assumptions rather than data. However, employing an ill-defined diffusion process to solve the IM problem may lead to suboptimal outcomes (Aral & Dhillon, 2018).
Uplift Modeling.  Another method for optimizing treatment allocation is UM (Devriendt et al., 2018). This approach involves estimating the ITE for each entity, ranking them according to these estimates, and selecting the top $k$ ranked entities with the highest expected ITE for treatment. An important limitation of this approach, however, is its assumption that entities are independent of each other. This assumption is violated in network settings where entities influence each other, such as marketing or vaccination campaigns. Recently, advanced causal inference methods have been proposed to cope with such network effects, known as spillover effects, by incorporating network information to estimate treatment effects (Forastiere et al., 2021; Jiang & Sun, 2022; Cai et al., 2023; Huang et al., 2023). However, even with the capability to estimate treatment effects in the presence of interference, determining the optimal treatment allocation has been largely overlooked. Ranking entities according to their estimated treatment effects is no longer a possibility because the decision to treat one entity may impact the decision whether or not to treat another entity.
Figure 1: Overview of OTAPI. In Step 1, a relational causal estimator is trained on historical data. It receives the treatment assignment $\mathbf{t}$ and the network $\mathcal{G}$, including the features $\mathbf{x}_{i}$ for each node $i$ (represented by the colors) as input. The estimator outputs the predicted outcomes $\hat{y}_{i}$ for each node. To train the model, these predictions are compared with the actual outcomes $y_{i}$. In Step 2, an optimization algorithm (e.g., a genetic algorithm) is used to find the optimal treatment allocation $\mathbf{t^{*}}$ for a potentially unseen network (e.g., a network that has changed over time). This optimization algorithm takes as input the new network and the set of feasible allocations, which depends on the budget $k$. The trained causal model is leveraged to estimate the total effect of treatment allocation, which is used as the objective function in the optimization algorithm.
OTAPI.  We propose OTAPI: Optimizing Treatment Allocation in the Presence of Interference, a novel method that combines the strengths of both UM and IM approaches to identify improved, data-driven, solutions to the IM problem. Our method consists of two steps. First, a relational causal estimator is trained using observational data, e.g., from historical marketing or vaccination campaigns. A relational causal estimator is an estimator that can leverage both individual and network information to estimate causal effects. This estimator allows estimating the total effect of any treatment allocation and controls for confounding. Second, the treatment effect estimates from this causal estimator are utilized to optimize the treatment allocation. Given the combinatorial nature of this optimization problem, this step presents a significant challenge. To address this challenge, treatment allocation is optimized by integrating treatment effect estimates into existing IM algorithms. An overview of OTAPI is depicted in Figure 1.
Contributions.  To the best of our knowledge, the challenge of optimally allocating treatment in the presence of interference is yet to be addressed. We conjecture that leveraging relational causal estimators will aid in finding the optimal treatment allocation. To this end, (1) we formalize the problem of finding the optimal treatment allocation in the presence of interference; (2) we present OTAPI, a new two-step method for optimizing treatment allocation in the presence of interference; and (3) we run extensive experiments on both synthetic and semi-synthetic data to empirically demonstrate that OTAPI finds better treatment allocations (in terms of the total effect of treatment allocation) than traditional methods used in UM and IM.
This article is structured as follows. In the following section, a comprehensive review of treatment effect estimation in the presence of interference and the IM problem is provided. Next, the problem setup and methodology are presented. Subsequently, we describe the experimental setup and analyze the results, before concluding this article and discussing the limitations and directions for future work.
2 Related work
2.1 Treatment effect estimation in the presence of interference
A substantial body of literature has focused on treatment effect estimation when entities are independent (Johansson et al., 2016; Yao et al., 2018; Devriendt et al., 2018). Within this literature, an important and extensively studied problem is confounding bias (Rosenbaum & Rubin, 1983; Shalit et al., 2017; Vairetti et al., 2024). This bias is associated with the treatment assignment mechanism, whereby treatment is not assigned randomly. Consequently, the distributions of the control and treatment groups will differ. In most applications, some form of mechanism or policy typically governs treatment assignment, leading to potential bias in treatment effect estimates. For instance, in marketing, a company might target a specific subgroup of the population. As it may be costly or even impossible to randomize the treatment assignment, (potentially biased) observational data has to be used.
Recent approaches construct balanced representations of the distributions of the control and treatment groups to mitigate confounding bias (Shalit et al., 2017; Johansson et al., 2016; Yao et al., 2018; Hassanpour & Greiner, 2019). In a network setting, where entities can influence each other, mitigating confounding bias becomes more challenging. The features and treatments of connected entities might influence the probability of receiving treatment, introducing confounding bias as well (Ma & Tresp, 2021; Jiang & Sun, 2022).
In estimating treatment effects in the presence of interference, a common approach involves defining an exposure mapping (Aronow & Samii, 2017), which generates a variable (or set of variables) based on the treatments of other relevant entities. Many methods assume this exposure mapping to be the percentage of directly connected entities, known as neighbors, that received treatment and use Graph Neural Networks (GNNs) (Kipf & Welling, 2016; Xu et al., 2018) to learn a model from observational data (Ma & Tresp, 2021; Jiang & Sun, 2022; Cai et al., 2023). Representation balancing techniques are employed to mitigate confounding bias. Some recent works propose to relax the assumption on the exposure mapping to allow for heterogeneous peer influence, where the spillover effect may vary depending on the features of connected entities (Huang et al., 2023; Adhikari & Zheleva, 2023). Another line of work focuses on leveraging network information, such as features of neighbors, to better control for hidden confounders. For example, a person’s community affiliations may help to represent their socioeconomic status. However, these methods typically ignore spillover effects, assuming that the treatment of one entity does not causally affect other entities (Guo et al., 2020, 2021).
2.2 Influence Maximization problem
The IM problem was initially introduced by Domingos & Richardson (2001), and later formalized as a combinatorial optimization problem by Kempe et al. (2003). A multitude of methods to tackle the IM problem rely on diffusion processes, which define how influence propagates through the network. A wide variety of diffusion processes have been proposed, each with its distinct assumptions about the dissemination of behavior or information (Goldenberg et al., 2001; Granovetter, 1978; Kempe et al., 2003; Ye et al., 2022). To identify the optimal allocation set, existing methods mostly employ a greedy algorithm proposed by Kempe et al. (2003). This algorithm uses Monte Carlo (MC) simulation based on the diffusion process to estimate the expected influence spread, i.e., the expected number of activated entities. However, its computational complexity may render it infeasible in terms of runtime, especially when networks become large. To this end, several methods have been developed to reduce the number of MC simulations, such as CELF (Leskovec et al., 2007) and CELF++ (Goyal et al., 2011). Other methods were developed to partition the network into communities, significantly reducing the runtime of an MC simulation (Wang et al., 2010; Bozorgi et al., 2017). Alternatively, some approaches ignore diffusion processes altogether, instead solely using the network structure. These methods employ, for instance, the degree centrality, which is defined as the number of connections of a node. IM algorithms based on this measure include the degree (Kempe et al., 2003) and single discount heuristics (Chen et al., 2009).
All the aforementioned methods rely on the network structure or make assumptions regarding the diffusion process in order to arrive at a solution. It has been shown, however, that incorporating data to estimate the spread of influence through the network yields more reliable solutions (Aral & Dhillon, 2018). In the IM literature, considerable work has been done to employ propagation traces to estimate the parameters of specific diffusion processes (Goyal et al., 2010; Saito et al., 2008; Bourigault et al., 2016; Tran & Zheleva, 2022). Propagation traces are data structures containing information about which entities were activated and when. Goyal et al. (2011) uses these traces to create a model that does not make any parametric assumptions about the diffusion process, but instead learns the expected influence spread directly from data. These methods, however, do not consider node features. The only exception is Tran & Zheleva (2022), in which node features are leveraged to estimate the parameters in the Linear Threshold model, a commonly used diffusion process. In contrast to these approaches, our proposed method leverages node features, does not make parametric assumptions regarding the diffusion process, and does not use propagation traces. Moreover, OTAPI allows for out-of-sample IM, as the network for which we try to find the optimal treatment allocation may be different from the network that was used for training the model. This is an important advantage of OTAPI, as networks – especially social networks – typically change over time. Ma & Tresp (2021) proposes a utility maximization approach for instance-level treatment assignment in the presence of interference using a policy network that leverages a relational causal estimator. The budget constraint is satisfied by integrating it into the loss function of the policy network and allowing some tolerance for violating it. In comparison, OTAPI directly optimizes treatment allocation for the network as a whole by leveraging algorithms that can support hard constraints.
3 Problem setup
Notation.  Consider an undirected network $\mathcal{G}=(\mathcal{N},\mathcal{E})$ where $\mathcal{N}$ represents the set of nodes and $\mathcal{E}$ denotes the set of edges between them. We define the set of neighbors of node $i$ as $\mathcal{N}_{i}$, which comprises nodes $j\in\mathcal{N}$ for which an edge $(i,j)$ exists. In this context, $(i,j)=(j,i)$. Additionally, we denote the set of nodes not in the neighborhood of $i$ as $\mathcal{N}_{-i}=\mathcal{N}\setminus(\mathcal{N}_{i}\cup\{i\})$. Each node $i$ possesses a feature vector $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$. We will use nodes and entities interchangeably, depending on the context. An entity is defined by an index $i$ and has a feature vector $\mathbf{X}_{i}$.
Let $Y_{i}\in\mathcal{Y}$ represent the outcome of entity $i$. Without loss of generality, $\mathcal{Y}=\{0,1\}$, e.g., buying or not buying the product. We define $T_{i}\in\mathcal{T}=\{0,1\}$ as the treatment given to entity $i$. The treatments of $i$’s neighbors $\mathcal{N}_{i}$ is given by $\mathbf{T}_{\mathcal{N}_{i}}=\{T_{j}\}_{j\in\mathcal{N}_{i}}$. Following (Forastiere et al., 2021; Ma & Tresp, 2021; Aronow & Samii, 2017), an exposure mapping, denoted by $Z:\mathbf{T}_{\mathcal{N}_{i}}\to Z_{i}$, is defined to summarize $\mathbf{T}_{\mathcal{N}_{i}}$. This function describes how the outcome of an entity is dependent on $\mathbf{T}_{\mathcal{N}_{i}}$. The domain $\mathcal{Z}$ of the exposure mapping $Z_{i}$ depends on the exact definition of the function $Z$ (Forastiere et al., 2021).
[…]
Figure 2: Graphical representation of the causal structure. Only two entities are included for simplicity. $Y_{i}$ is influenced by its own features $\mathbf{X}_{i}$, the features of its neighbors $\mathbf{X}_{j}$, its own treatment $T_{i}$, and the treatment of its neighbors $T_{j}$. Moreover, $T_{i}$ is influenced by both $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$
Assumptions.  We assume that an entity’s own treatment $T_{i}$, the treatments of its neighbors $\mathbf{T}_{\mathcal{N}_{i}}$, an entity’s features $\mathbf{X_{i}}$ and the features of its neighbors $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$ causally influence its outcome $Y_{i}$. This causal structure is depicted in Figure 2. Note that we do not consider contagion effects, meaning the outcomes of different entities do not causally affect each other (Ogburn & VanderWeele, 2014). To summarize $\mathbf{T}_{\mathcal{N}_{i}}$ we use the mapping $Z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}T_{j}}{|\mathcal{N}_{i}|}$ (Forastiere et al., 2021; Jiang & Sun, 2022). Note that more complex exposure mappings (e.g., multi-dimensional, based on node features, etc.), can also be defined/learned (Cai et al., 2023; Adhikari & Zheleva, 2023; Aronow & Samii, 2017). We assume that the decision to treat is taken simultaneously for all entities. Consequently, both $Z_{i}$ and $T_{i}$ are part of the intervention on $i$.
Traditionally, each entity has two potential outcomes: $Y_{i}(T_{i}=0)$ and $Y_{i}(T_{i}=1)$. However, in the presence of interference, the traditional consistency assumption, which states that there are no multiple versions of treatment, is violated. Here, depending on the treatment of $i$’s neighbors, $i$’s potential outcomes might differ. To address this, the consistency assumption is relaxed by employing the exposure mapping $Z_{i}$ (Forastiere et al., 2021; Jiang & Sun, 2022):
Consistency. If $T_{i}=t$ and $Z_{i}=z$, then $Y_{i}=Y_{i}(t,z)$.
Overlap and strong ignorability. Moreover, we introduce the overlap and strong ignorability assumptions in the presence of interference (Forastiere et al., 2021; Jiang & Sun, 2022): $\exists\ \delta\in(0,1)$ such that $\delta<p(T_{i}=1|\mathbf{X}_{i}=\mathbf{x}_{i},\mathbf{X}_{\mathcal{N}_{i}}=\mathbf{x}_{\mathcal{N}_{i}})<1-\delta$ and $Y_{i}(T_{i}=t,Z_{i}=z)\perp\!\!\!\perp$ $T_{i},Z_{i}$ $|\mathbf{X}_{i},\mathbf{X}_{\mathcal{N}_{i}}$, $\forall t\in\mathcal{T}$, $z\in\mathcal{Z}$, $\mathbf{X}_{i}\in\mathcal{X}$ and $\mathbf{X}_{\mathcal{N}_{i}}=\{\mathbf{X}_{j}\}_{j\in\mathcal{N}_{i}}$. However, the traditional ignorability assumption no longer holds: $Y_{i}(T_{i}=t)\centernot{\perp\!\!\!\perp}$ $T_{i}$ $|\mathbf{X}_{i}$ $\forall t\in\mathcal{T}$ and $\mathbf{X}_{i}\in\mathcal{X}$.
Treatment effect definitions.  Based on the introduced assumptions and notation, the marginal individual treatment effect (MITE) is defined as:
$\tau_{i}(z)=\mathbb{E}\bigl{[}Y_{i}(1,z)-Y_{i}(0,z)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]},$ (1)
with $Y_{i}(t,z)=Y_{i}(T_{i}=t,Z_{i}=z)$. Equation 1 shows that MITE is dependent on the value of $Z_{i}$. Note that the MITE differs from the traditional individual treatment effect (ITE) as the latter does not take into account graph information and is traditionally defined as follows:
$\tau_{i,trad}=\mathbb{E}\bigl{[}Y_{i}(T_{i}=1)-Y_{i}(T_{i}=0)|\mathbf{x}_{i}\bigr{]}.$ (2)
The above equation is ill-defined in the presence of interference since $Y_{i}$ not only depends on $i$’s treatment and features but also on the treatments and features of $i$’s neighbors. We can adapt the traditional definition of the ITE to a network context as follows:
$\tau_{i,trad}=\mathbb{E}_{(\mathbf{X}_{\mathcal{N}_{i}},Z_{i})}\bigl{[}\mathbb{E}\left[Y_{i}(T_{i}=1,Z_{i}=z)-Y_{i}(T_{i}=0,Z_{i}=z)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\right]\bigr{]}.$ (3)
In this definition, we marginalize out $\mathbf{X}_{\mathcal{N}_{i}}$ and $Z_{i}$. This is what classical causal inference techniques would estimate in the presence of interference.
The spillover effect is defined as:
$\delta_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(t,0)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (4)
The spillover effect is the effect of the treatments of $i$’s neighbors on $i$, and not the effect of the treatment of $i$ on its neighbors. Furthermore, the spillover effect is dependent on the value of $T_{i}=t$.
Finally, we can also define the individual total treatment effect (ITTE) as:
$\omega_{i}(t,z)=\mathbb{E}\bigl{[}Y_{i}(t,z)-Y_{i}(0,0)|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{]}.$ (5)
The ITTE quantifies the total effect of the full treatment allocation on an entity. It is noteworthy that the ITTE can be rewritten in terms of MITE and spillover effect:
$\omega_{i}(t,z)=\begin{cases}\delta_{i}(0,z)&\text{if }t=0,\\ \tau_{i}(0)+\delta_{i}(1,z)=\delta_{i}(0,z)+\tau_{i}(z)&\text{if }t=1.\end{cases}$ (6)
Objective.  We aim to find an optimal treatment allocation $\mathbf{t^{*}}=\{t_{i}\}_{i\in\mathcal{N}}$ that maximizes the total treatment effect (TTE). The TTE is defined as the sum of ITTE $\sum_{i\in\mathcal{N}}\omega_{i}(t_{i},z_{i})$ and quantifies the total effect of treatment allocation on all entities. The problem can now be formalized as:
$\mathbf{t^{*}}=\underset{\mathbf{t}}{\arg\max}\sum_{i\in\mathcal{N}}\omega_{i}(t_{i},z_{i})\text{ s.t. }\sum_{i\in\mathcal{N}}t_{i}\leq k.$ (7)
Both $t_{i}$ and $z_{i}$ are decision variables and can be derived from $\mathbf{t}$. Now, if we have a model that can estimate the ITTE for a given $\mathbf{t}$, we can use (the sum of) these estimates as our objective function to find an optimal treatment allocation $\mathbf{t^{*}}$.
4 OTAPI
To clearly position OTAPI, we first compare it to two alternative treatment allocation approaches: Uplift Modeling (UM) and classic Influence Maximization (IM).
In UM, the focus lies on finding the entities with the highest expected ITE. The goal in UM is essentially the same as (7). To maximize TTE, entities are ranked according to estimated ITE, and the top $k$ entities are given treatment. Alternatively, integrated approaches exist that rank entities directly without estimating ITE first (Devriendt et al., 2020). Nevertheless, these approaches are not tailored to scenarios involving network interference because they rely solely on individual information. When network interference plays a significant role, ITE estimates may be biased, potentially resulting in suboptimal treatment allocation, where entities with substantial network effects might be overlooked and connected entities might all be treated while treating just one of them might have a similar effect in terms of TTE (Forastiere et al., 2021). Moreover, each treatment allocation vector $\mathbf{t}$ may result in a different ITTE $\omega_{i}$ in the presence of interference. Therefore, simply ranking entities according to ITTE no longer works because an entity’s ITTE depends on the set of other entities that receive treatment.
Classic IM methods focus on modeling network effects, often using diffusion processes to simulate interference dynamics. These require computationally expensive MC simulations to estimate TTE. These TTE estimates lack any form of individualized treatment effect estimation and are commonly based on assumptions regarding the applicable diffusion process. By leveraging heuristics, the TTE is maximized without having to run simulations for every possible treatment allocation.
OTAPI addresses these limitations by combining the strengths of these two methods. It consists of two components: (i) a relational causal estimator that considers both individualized susceptibility to treatment and interference to estimate ITTE (Section 4.1), and (ii) an allocation algorithm that utilizes the predictions made by the trained causal model to allocate treatments optimally (Section 4.2).
4.1 Estimating treatment effects in the presence of interference
In order to optimally allocate treatment, accurate estimates of the ITTE $\omega_{i}(t_{i},z_{i})$ are required. To estimate unbiased treatment effects, we employ causal machine learning. OTAPI is agnostic to the estimator employed, but its performance directly depends on the accuracy of the estimates.
Given the assumptions in Section 3, we utilize a causal estimator $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$ which provides unbiased estimates for the potential outcomes $\mathbb{E}\bigl{(}Y_{i}(T_{i}=t_{i},Z_{i}=z_{i})|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}}\bigr{)}$. Since an entity can have an arbitrary number of neighbors, a GNN is employed to learn an aggregation of the neighbors’ features. Jiang & Sun (2022) shows that if the distributions $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$, the classic propensity score, and $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$, the peer exposure score, are biased, the estimates for the treatment effects will also be biased. Given that treatments are typically not allocated randomly in observational data due to confounding, controlling for these biased distributions is necessary.
To address the challenge of confounding bias in the presence of interference, representation balancing techniques can be applied (Ma & Tresp, 2021; Jiang & Sun, 2022; Shalit et al., 2017). This leads to more accurate estimates of the ITTE. In the experiments presented in Section 5, we adopt NetEst (Jiang & Sun, 2022), a GNN-based causal estimator that employs an adversarial approach to balance representations. Nevertheless, alternative relational causal estimators could also be employed.
4.2 Allocating treatments optimally
Assuming we can compute unbiased ITTE estimates for all entities as $\hat{\omega}_{i}(t_{i},z_{i})=\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i},z_{i})$- $\mathcal{M}(\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},0,0)$, we can plug these estimates into the objective (Equation 7). As already explained, the strategy of ranking entities according to estimated ITTE $\hat{\omega}_{i}$ no longer works. To address this, we rely on heuristics developed in the IM literature and plug data-driven estimates of TTE into these algorithms. Generally, these heuristics maximize TTE by iteratively improving a (set of) solution(s). Some examples include the greedy algorithm (Kempe et al., 2003), the genetic algorithm (Bucur & Iacca, 2016), and simulated annealing (Jiang et al., 2011). OTAPI can use any of these algorithms to optimize treatment allocation.
5 Experiments
5.1 Experimental setup
5.1.1 Data
To properly evaluate treatment allocation methods in the presence of interference, the ground truth TTE is required. However, due to the fundamental problem of causal inference (Holland, 1986), the TTE is unobserved. Therefore, in causal inference, (semi-)synthetic data is often used (see, e.g., Vanderschueren et al., 2023; Berrevoets et al., 2020). Hence, we generate synthetic and semi-synthetic data according to the causal diagram in Figure 2.
Synthetic data.  For each synthetic dataset, 3 separate networks (train, validation, and test) are generated according to the Barabási-Albert random network model (Barabási & Albert, 1999). Each network consists of 5,000 nodes. The hyperparameter which determines the number of edges attached from each new node is set to $m=2$. The data-generating process (DGP) is inspired by Jiang & Sun (2022).
For each node, a feature vector of $10$ features is randomly sampled according to a standard normal distribution:
$\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
with $\mathbf{I}$ the identity covariance matrix. We first generate the parameters that will be used later on in the DGP:
[…]
Where Unif is the uniform distribution. These parameters are used to quantify the effect of $\mathbf{X_{i}}$ on ${T_{i}}$, $\mathbf{X_{i}}$ on $Y_{i}$, and the effect of $\mathbf{T_{i}}$ on $Y_{i}$ respectively. The treatment $t_{i}$ is generated as follows:
$&t_{i}\sim\text{Bernoulli}(p_{t_{i}});&p_{t_{i}}=\beta_{XT}\cdot p_{i}+\beta_{{\mathcal{N}}T}\cdot p_{\mathcal{N}_{i}},\\ \end{aligned}$
with $p_{i}=\sigma(\mathbf{w}^{XT}\cdot\mathbf{x}_{i})$ and $p_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}p_{j}}{|\mathcal{N}_{i}|}$. Here, $\sigma$ is the sigmoid function. Note that confounding is introduced by generating propensity scores $p_{i}$ and $p_{\mathcal{N}_{i}}$ based on $\mathbf{x}_{i}$ and $\mathbf{x}_{\mathcal{N}_{i}}$, respectively. Now, we describe how the outcome $y_{i}$ is generated:
[…]
where $h_{i}=\sigma(\mathbf{w}^{TY}\cdot\mathbf{x}_{i})+b_{TY}$, $u_{i}=\sigma(\mathbf{w}^{XY}\cdot\mathbf{x}_{i})$, and $u_{\mathcal{N}_{i}}=\frac{\sum_{j\in\mathcal{N}_{i}}u_{j}}{|\mathcal{N}_{i}|}$. Recall that $z_{i}$ is the ratio of $i$’s neighbors that received treatment $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. $h_{i}$ is used to make the treatment effect heterogeneous, i.e., dependent on the features. Note that the individual effect is 100% correlated with the spillover effect. This choice is motivated by the fact that if someone is more/less susceptible to treatment, they will most likely also be more/less susceptible to the treatment of neighbors. The following setting for the parameters was used for all experiments:
[…]
The intuition behind these parameters is that the individual features and treatments will always be more impactful than those of neighbors. $\beta_{spillover}$ is varied between 0 and 0.7 across experiments.
Semi-synthetic data.   Following Jiang & Sun (2022) and Guo et al. (2020), we use two real social network datasets: BlogCatalog (BC) and Flickr. We follow the same procedure as Jiang & Sun (2022) to generate the networks and features from the Flickr and BC datasets. Some summary statistics of these datasets are given in Table 1.
Flickr BC
Train Validation Test Train Validation Test
Nodes 2,482 2,461 2,358 1,716 1,696 1,784
Edges 46,268 14,419 23,529 17,937 25,408 14,702
Features 12,047 12,047 12,047 8,189 8,189 8,189
Table 1: Flickr and BC datasets
These networks are split using METIS (Karypis & Kumar, 1998) which results in three separate networks: a train, validation, and test set. We directly use the same split from Jiang & Sun (2022). Given the large number of (sparse) features, we leverage LDA (Blei et al., 2003) to reduce this number to 10 for both datasets. Now, the same DGP as in the synthetic datasets is used. For the semi-synthetic datasets, $\beta_{spillover}=0.3$. All other parameters are the same.
$\mathbf{x}_{\mathcal{N}_{i}}$$\mathbf{x}_{i}$$\mathbf{\phi}_{i}$$t_{i}$$z_{i}$$e_{\phi}$$d_{T}$$d_{Z}$$p_{Y}$$\hat{y}_{i}$$\hat{z}_{i}$$\hat{t}_{i}$
Figure 3: NetEst architecture
5.1.2 Causal estimator
As already mentioned in Section 4, we will use a modified version of NetEst (Jiang & Sun, 2022) to estimate ITTEs. We opt for NetEst because of its strong empirical performance, solid theoretical foundation, and the availability of its source code. The architecture is shown in Figure 3. Inputs are given in blue  , (intermediate) outputs in yellow  , and the neural network layers in green  . NetEst uses an adversarial training approach to balance the representations $\mathbf{\phi}_{i}$ (Berrevoets et al., 2020). This is achieved by using two discriminators $d_{T}$ and $d_{Z}$ that predict the treatment $t_{i}$ and the peer exposure $z_{i}$, respectively. To balance the representations, the distributions $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$ and $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$ are uniformed. More specifically, there should be no information in the hidden representation $\phi_{i}$ that allows to predict $t_{i}$ or $z_{i}$. To calculate $\phi_{i}$, a Graph Convolutional Network (GCN) (Kipf & Welling, 2016) and an MLP $e_{\phi}$ are used. $\phi_{i}$ is fed into the predictor $p_{Y}$ to make a prediction for the potential outcome ${y}_{i}$. The discriminators $d_{Z}$ and $d_{T}$, and the predictor $p_{Y}$ are MLPs with ReLU non-linearities that consist of 3 layers with each hidden layer consisting of 16 hidden nodes. $e_{\phi}$ has one layer with a ReLU non-linearity. The GCN consists of one convolutional layer (Jiang & Sun, 2022).
Every epoch consists of two steps. First, the discriminators are trained to accurately predict $T_{i}$ and $Z_{i}$. To this end, Binary Cross Entropy (BCE) and Mean Squared Error (MSE) are used to respectively update the discriminators $d_{T}$ and $d_{Z}$. In the second step, the discriminators are frozen. Then, regularization losses are calculated based on the discriminators as follows:
$\mathcal{L}_{uT}=\frac{1}{N}\sum_{i\in\mathcal{N}}(\hat{t_{i}}-0.5)^{2}\text{ and }\mathcal{L}_{uZ}=\frac{1}{N}\sum_{i\in\mathcal{N}}(\hat{z_{i}}-c_{i})^{2}.$ (8)
These regularization losses are employed to uniform the distributions. Since $T_{i}$ is binary, we can use $0.5$ as the uniform distribution. For $Z_{i}$, $c_{i}$ is sampled uniformly at random from $\left[0,1\right]$ for each $i$ in each iteration. The outcome loss $\mathcal{L}_{Y}$ compares the prediction $\hat{y}_{i}$ with the actual outcomes $y_{i}$ using Binary Cross Entropy (BCE). $\mathcal{L}_{Y}$ is used to update the estimator $e_{Y}$. Finally, the combined loss is calculated:
$\mathcal{L}_{comb}=\mathcal{L}_{Y}+\alpha\mathcal{L}_{uT}+\gamma\mathcal{L}_{uZ}.$
This loss is used to update the GCN and the encoder $e_{\phi}$ (Jiang & Sun, 2022).
The main modifications made to the original NetEst architecture are the change to a binary outcome $Y_{i}$ and the addition of the encoder $e_{\phi}$. The latter improves performance by separating $\mathbf{x}_{i}$ from $\mathbf{x}_{\mathcal{N}_{i}}$. Following Jiang & Sun (2022) we set $\alpha$ and $\gamma$ to 0.5. The Adam optimizer (Kingma & Ba, 2015) is employed to train the network. Furthermore, the neural network is trained using full-batch training. Code provided by Jiang & Sun (2022) was used as a starting point for both the DGP and NetEst implementation.
5.1.3 Optimization algorithms
To maximize TTE we use two heuristics: the greedy algorithm (OTAPI-GR) (Kempe et al., 2003) and the genetic algorithm (OTAPI-GA) (Bucur & Iacca, 2016). The greedy algorithm is an iterative approach where the allocation set is sequentially expanded until the budget $k$ is reached. In each iteration, the marginal gain in TTE associated with the addition of each remaining node to the allocation set is calculated. This marginal gain can be estimated as $\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i}^{\prime},z_{i}^{\prime})-\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i},z_{i})$ where $\mathbf{t}$ is the treatment vector from the previous iteration, and $\mathbf{t^{\prime}}$ is $\mathbf{t}$ with an additional node $j$ receiving treatment. The treatment vector $\mathbf{t^{\prime}}$ with the highest estimated TTE is chosen for that iteration.
The genetic algorithm (GA) iteratively improves a population of solutions, represented by treatment allocation vectors $\mathbf{t}$, by applying genetic operators such as crossover and mutation. First, a fitness function has to be defined. We employ predicted TTE. To incorporate the budget constraint, the fitness function returns zero when this constraint is violated. An initial set of solutions has to be generated. To converge more quickly, and to get better solutions, the degree (DEG) and single discount (SD) heuristic solutions are added to the initial population. The other solutions in the initial population are randomly generated. Thereafter, the best solutions are selected to be parents to create new solutions (children). These solutions differ from their parents by using crossover (a child is made up of the parts of its parents) and mutation (randomly changing small parts of the child). Thanks to this randomness, escaping from local optima is made possible. Finally, – to not lose any really good solutions – the best solutions from the previous iteration are kept unchanged and the next iteration starts (Gad, 2023; Bucur & Iacca, 2016). The genetic algorithm was implemented using the PyGAD library (Gad, 2023). The hyperparameters were tuned using the predicted TTE on the synthetic data with $\beta_{spillover}=0.3$. The hyperparameters are shown in Table 2. The number of generations is dependent on the budget $k$ because the algorithm converges faster for lower $k$.
Parameter Value
Num. generations $\begin{cases}37\cdot k+300&\text{if }k\leq 100\\ 5,000&\text{if }k>100\end{cases}$
Population size $40$
Num. elites $5$
Num. parents mating $15$
Num. genes mutated $1$
Crossover type uniform
Table 2: Hyperparameters GA
5.1.4 Methods for comparison
We implemented the following methods to compare OTAPI with:
  • Degree (DEG): pick the $k$ nodes with highest degree.
  • Single Discount (SD): stepwise degree as described by Chen et al. (2009). In each iteration, the node with the highest degree is added to the allocation set. Once a node is selected, all its edges are deleted and the degree is recalculated. This process stops when the budget $k$ is reached.
  • CELF: IM approach. A variation of the greedy algorithm that reduces the required amount of MC simulations per iteration by leveraging the submodularity property that some diffusion processes exhibit (Leskovec et al., 2007). We employ the Independent Cascade (IC) model with probability $p=0.01$ as the diffusion process because of its popularity within the IM literature (Kempe et al., 2003). We simulate the diffusion process $1000$ times per step to get an accurate estimate of the TTE.
  • TARNet: UM approach. TARNet (Shalit et al., 2017), a commonly used causal estimator, is used for ITE estimation. TARNet is trained on individual data only, not leveraging any network information. The $k$ entities with the highest ITE estimates are allocated treatment. We use the implementation for TARNet (Shalit et al., 2017) from Jiang & Sun (2022) and largely follow their setup. The model is trained using full-batch training with BCE loss and employs the Adam optimizer Kingma & Ba (2015). The number of hidden nodes in each layer is 16.
  • Upper Bound (UB): To get an idea of the upper bound on the evaluation metrics, we implement a method that leverages the true DGP, which simulates the performance of a perfect model $\mathcal{M}$. Since the allocation problem is – even with a perfect model – still a combinatorial problem, a heuristic has to be used to solve it. To this end, we use the greedy heuristic. Note that this means that the true upper bound may be better than obtained with this heuristic.
5.1.5 Evaluation metrics
We define Liftup in TTE for method $m$ and budget $k$ as
$Liftup_{m}(k)=\frac{TTE_{m}(k)}{TTE_{random}(k)}.$ (9)
This metric captures the relative increase in TTE of a method compared to randomly assigning an equal amount of treatments. Since there are many possible random assignments, we randomly sample 100 possible assignments for a budget $k$ and calculate the average TTE to estimate $TTE_{random}(k)$. 100 was chosen because it gave stable estimates for all $k$. Moreover, to get a more complete understanding of the results, the relative increase in sum of expected outcomes (RISEO) is defined as:
$RISEO_{m}(k)=\frac{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,m},z_{i,m})\bigr{]}}{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,random},z_{i,random})\bigr{]}}.$ (10)
Here, $t_{i,m}$ and $z_{i,m}$ result from the treatment allocation by method $m$. This metric quantifies the relative increase in the sum of expected outcomes (e.g., total sales) compared to random treatment allocation. We use the same strategy to estimate the value of random treatment allocation as in Liftup.
5.1.6 Hyperparameter selection
We use grid search to find the best hyperparameters for TARNet and NetEst for each setup. To select the best model we use the validation Binary Cross Entropy (BCE) to compare the factual outcomes $y_{i}$ to the predicted outcomes $\hat{y}_{i}$. However, the model that is best at predicting the potential outcomes, is not necessarily the best at predicting causal effects. Although model selection in causal inference is still an unresolved issue, better approaches may exist (Curth & Van Der Schaar, 2023). However, given that these approaches have not yet been tested in a network context, we leave this for future research. The hyperparameter ranges for NetEst and TARNet are shown in Tables 3 and 4. The source code for all our experiments can be found in this repository: https://github.com/daan-caljon/OTAPI.
Parameter Range
Learning rate $\{5\cdot 10^{-3},1\cdot 10^{-3},5\cdot 10^{-4}\}$
Epochs $\{200,500,800\}$
Table 3: Hyperparameters NetEst
Parameter Range
Learning rate $\{5\cdot 10^{-3},1\cdot 10^{-3},5\cdot 10^{-4}\}$
Epochs $\{500,800\}$
Num. representation layers $\{1,2\}$
Num. hypothesis layers $\{1,2\}$
Table 4: Hyperparameters TARNet
5.2 Results and discussion
5.2.1 Influence of budget $k$
In LABEL:fig:semi-synthetic, we show the results for the test sets of the two semi-synthetic datasets BC and Flickr. We plot Liftup and RISEO against the budget $k$ and observe that OTAPI-GR and OTAPI-GA outperform (or have at least the same performance as) all other methods for all $k$. Furthermore, their performance is close to the Upper Bound, especially for the Flickr dataset. Note that OTAPI-GR and OTAPI-GA have an almost identical performance (making it hard to see the curves of OTAPI-GR in the figures). Moreover, "network-aware" methods (i.e., all methods except for TARNet) outperform the UM approach for relatively small budgets $k$. However, once a certain budget is surpassed, the TARNet method outperforms CELF, DEG, and SD for both datasets. This happens already for a small budget $k$ for the BC dataset. This pattern is expected since the degree distribution follows a power law (Barabási & Albert, 1999). This means that a relatively low number of nodes have a high degree. Intuitively, we would expect that treating these nodes should have a large impact on the network as a whole, which is confirmed by these results. This is further explored in Section 5.2.3. The UM approach (TARNet) fails to identify these high-impact entities. As $k$ increases, identifying nodes with a large marginal individual effect becomes more important than the spillover effects. OTAPI can capture both the spillover and marginal individual effects, which explains why it performs well for all $k$. Note that the peaks of the Liftup and RISEO curves have no direct implications in terms of decision-making. The optimal $k$ depends on the relative costs and benefits of the treatments and their effects, respectively (Verbeke et al., 2020).
5.2.2 Influence of the magnitude of the spillover effect
An important parameter in the DGP is $\beta_{spillover}$. This parameter determines the average magnitude of the spillover effect. To examine the impact of this parameter on the relative performance of the different methods, we generate synthetic data consisting of 5,000 nodes and gradually increase $\beta_{spillover}$. We show the Liftup at different values of $k$ on the test set for all methods for $k=1\%$ and $k=5\%$ and for all non-greedy methods for higher budgets $k$ in LABEL:fig:spillover_effect. The greedy methods were not run for higher budgets due to a high runtime. From the results of this experiment, we can draw four interesting conclusions. (1) OTAPI-GR and OTAPI-GA outperform all benchmarks for all values of $\beta_{spillover}$. (2) The TARNet method performs worse as the spillover effect increases and has the same performance as OTAPI for $\beta_{spillover}=0$. The latter is expected since this is the classic UM setting with no interference. (3) The relatively simple DEG and SD heuristics exhibit performance more closely aligned with OTAPI as the average spillover effect increases. (4) As already concluded from LABEL:fig:semi-synthetic, we observe that TARNet becomes more competitive as $k$ increases. Furthermore, DEG and SD perform worse as $k$ increases and do not even outperform TARNet when $k=60\%$ and $k=80\%$ for a large $\beta_{spillover}$. The fact that OTAPI-GR becomes challenging in terms of runtime as $k$ increases also highlights the advantage of OTAPI-GA compared to OTAPI-GR. Given a certain budget $k$, OTAPI-GA only needs to calculate the best solution for that value of $k$. OTAPI-GR, on the other hand, has to calculate all solutions for smaller $k$ first.
5.2.3 Impact of the degree distribution
All networks investigated in this paper possess the scale-free property (Barabási & Albert, 1999). This means that the distribution of the degree follows a power law. As shown in LABEL:fig:degree_effect, a small number of nodes has a very high degree. Consequently, they will most likely have a large network effect. This may explain why the degree and single discount heuristics perform quite well for small $k$. To further investigate this hypothesis, we simulate a small dataset of 2,000 nodes using the Watts-Strogatz (Watts & Strogatz, 1998) random network model. Networks that follow this model do not possess the scale-free property. The degree distribution for the simulated data using the Watt-Strogatz model is shown in Figure LABEL:fig:Degree_WS. To make a fair comparison to the Barabási–Albert random networks, the parameters of the Watts-Strogatz random network were chosen to have the same average degree (4).
We set $\beta_{spillover}=0.3$ and run the experiment for all $k$. The results are shown in Figures LABEL:fig:liftup_WS and LABEL:fig:influence_WS. These differ significantly from previous results. TARNet performs very well even though there is a network effect. It outperforms the methods that only look at network structure (DEG, SD, and CELF). Moreover, these methods are not better than randomly allocating treatments. Although the results for the baseline methods are different, the results for OTAPI remain the same. It still outperforms (or has the same performance) as the best baseline.
5.2.4 Allocation similarity between methods
To gain more insight into the similarities and differences between methods, we calculate a similarity matrix between the solutions for the Flickr test set at $k=5\%$, $k=20\%$, and $k=40\%$. Similarity is defined as the number of selected nodes that two methods have in common divided by the budget $k$. The similarity matrices are shown in LABEL:tab:sim_matrix. As expected, solutions with a larger Liftup (see LABEL:fig:Flickr_Liftup) are more similar to the UB solution. Furthermore, we observe that OTAPI-GR and OTAPI-GA yield an almost identical solution. The high similarity explains why the Liftup for both methods is almost identical. This similarity between the two methods persists for higher $k$. Furthermore, the ability to leverage network structure plays a significant role in finding solutions with high similarity to UB. This pattern changes as $k$ increases, then the similarity between UB and TAR, and OTAPI and TAR increases, while the similarity between UB and DEG, and UB and SD decreases. This is because, at some point when increasing $k$, the marginal individual effects gradually become more important than the spillover effects. If this analysis were to be extended to higher $k$, it would become increasingly difficult to extract insightful conclusions from the similarity matrix as a random solution would already have an expected similarity of $k\%$ with each solution. Consequently, $k=40\%$ is the highest budget we investigate using the similarity matrix.
6 Conclusion
We present OTAPI, a novel method to optimize treatment allocation in the presence of network interference. Currently, two alternative approaches, IM and UM, exist. OTAPI fills the gap between these methods by combining their strengths. Our method consists of two steps. First, a relational causal estimator is trained to accurately estimate treatment effects in the presence of interference. Second, because the problem of finding the optimal treatment allocation in a network setting is combinatorial in nature, optimization algorithms are leveraged. These algorithms intelligently employ the relational causal estimator for treatment allocation.
We have empirically demonstrated that OTAPI outperforms alternative treatment allocation methods. Moreover, the relative performance of our method is robust to both the amount of available treatments $k$ and the magnitude of the spillover effects. By contrast, the relative performance of IM and UM approaches is highly dependent on these parameters.
A limitation of our work is that the assumed causal structure might not fully capture reality, as it does not account for potential contagion effects, which means that the outcome of an entity at time $t$ can influence the outcome of another entity at time $t+1$ (Ogburn & VanderWeele, 2014). Moreover, we have followed existing literature (Forastiere et al., 2021; Jiang & Sun, 2022) by assuming that the exposure mapping $Z$ is defined as the ratio of neighbors treated. This assumption will likely be violated in practice. Consequently, the causal estimator that was used in this paper may become quite inaccurate. However, since OTAPI is agnostic to the relational causal estimator that is used, other estimators that do not need this assumption (e.g., Huang et al., 2023; Adhikari & Zheleva, 2023) can be seamlessly incorporated. Nevertheless, OTAPI’s performance using more complex relational causal estimators should be validated in future work. Furthermore, the estimation errors from the relational causal estimator might have an impact on the treatment allocation optimization in the second step of OTAPI. Therefore, an interesting avenue for future research is to investigate whether an end-to-end approach, where the treatment allocation is directly learned from the input data, is feasible for this problem. Finally, also incorporating potential costs and benefits of treatments and outcomes provides an interesting direction for future research to determine the optimal number of treatments $k^{*}$ in a data-driven manner (Verbeke et al., 2020, 2023).
7 Acknowledgements
Wouter Verbeke and Daan Caljon thankfully acknowledge the financial support from Research Foundation - Flanders (project G038723N). Jente Van Belle gratefully acknowledges the financial support from Research Foundation – Flanders (project 12AZX24N).